# Exploring Neuron Interactions and Emergence in LLMs: From the Multifractal Analysis Perspective

### v4

@@ Abstract
Prior studies on the emergence in large models have primarily focused on how the functional capabilities of large language models (LLMs) scale with model size. Our research, however, transcends this traditional paradigm, aiming to deepen our understanding of the emergence within LLMs by placing a special emphasis not just on the model size but more significantly on the complex behavior of neuron interactions during the training process. By introducing the concepts of “self-organization” and “multifractal analysis,” we explore how neuron interactions dynamically evolve during training, leading to “emergence,” mirroring the phenomenon in natural systems where simple micro-level interactions give rise to complex macro-level behaviors. To quantitatively analyze the continuously evolving interactions among neurons in large models during training, we propose the Neuron-based Multifractal Analysis (NeuroMFA). Utilizing NeuroMFA, we conduct a comprehensive examination of the emergent behavior in LLMs through the lens of both model size and training process, paving new avenues for research into the emergence in large models.
Large language model, Multifractal analysis, Interpretability, Emergence
@@ 1 Introduction
The advent of large models triggered a paradigm shift in artificial intelligence (AI), demonstrating unparalleled capabilities across many domains (e.g., natural language processing Devlin et al.; Raffel et al.; Brown et al.; Chowdhery et al.; He et al.). Recent studies observed that, compared to smaller scale neural networks (NNs), the large language models (LLMs) can exhibit advanced cognitive functions and a heightened level of comprehension (Bubeck et al., 2023). This phenomenon has been encapsulated in a scaling law of performance, indicating a linear correlation between the model’s reducible loss and its size on a logarithmic scale (Kaplan et al., 2020; Hu et al., 2023a, b). Moreover, a sudden leap of model performance was observed when the model scale reaches a large level, being defined as the “emergence” phenomenon (Srivastava et al., 2022; Wei et al., 2022a).
Figure 1: Difference between traditional methods and our new method. We present a different perspective of how the model gets emergence from the training steps scale.
@@
Figure 2: Improvement of pythia responses with increasing training iterations, as indicated by the progression from incoherent to coherent answers across diverse categories (yellow for open-ended questions, blue for questions with standard answers, and red for semi-open-ended questions).
Empirical studies on analyzing the LLMs’ test performance identified power-law relationships but did not provide the intrinsic characteristics and rationale behind the emergence behavior. Previous research on the emergence in LLMs (Wei et al., 2022a; Fu et al., 2023) focused on model scale (e.g., number of parameters, model depth), but ignored the evolution of model performance during training. The iterative training process is essential for LLM development, demonstrating their evolving capabilities over time, as illustrated in Fig. 2. Large models often require more comprehensive training to achieve heightened abilities. Besides, traditional metrics (Paperno et al., 2016; Sakaguchi et al., 2021; Bisk et al., 2020) only evaluate the LLM’s output performance. Moreover, LLMs exhibit variations in their internal structures and dynamic characteristics during the training process (Ding et al., 2023). Although previous work may observe the external manifestations of emergence, they overlook the complex dynamics and internal interactions during the model training process, which are key to understanding the emergence.
Our study, therefore, extends beyond observable outcomes to explore the intricate mechanisms driving emergence in neural networks. We suggest that emergence results from the dynamic interplay and evolving complexity of neuron interactions during training, leading to sophisticated order / patterns. This mirrors the “self-organization” seen in natural systems, where simple rules at the micro-level lead to complex macro-level behaviors (see Appendix A.3). To decipher how collective neuronal interactions contribute to the self-organization and emergence phenomena during LLM’s training process, we construct a network representation of the neuronal interactions in LLMs and propose the Neuron-based Multifractal Analysis (NeuroMFA) (see Methodology 4.2) to characterize the pattern formation during these evolving interactions. NeuroMFA can learn the multiscale patterns that result from complex interactions among neurons and helps us unravel the subtle dynamics, employing a specialized weighted box covering method tailored for models’ directed weighted structure. Specifically, NeuroMFA enables precise measurement of neuron-based fractal properties and introduces the distortion factor to distinguish these structures with multiple fractal properties accounting for common and rare patterns. This multifaceted approach not only assesses the multifractal nature of neuron interactions but also deepens our understanding of the model’s training progression and the emergence of intelligence within LLMs.
In summary, the contributions of our work are three-fold: (i) We present a collective systems-inspired framework that represents LLMs as neuron interaction networks (NINs), enabling inter-neuron connectivity analysis; (ii) we introduce the neuron-based multifractal analysis (NeuroMFA) to quantify the regularity and heterogeneity of NINs, allowing us to understand neuron interactions and collective behaviors from a self-organization perspective; (iii) based on NIN and NeuroMFA, we propose a metric to measure the degree of emergence in LLMs from the perspective of neuron self-organization during the training process, opening a new direction for the study of emergence in large models.
The structure of this paper is as follows: Section 2 and 3 introduce the basic idea of emergence and self-organization, along with related work. Section 4 summarizes our NeuroMFA framework for quantifying the emergence phenomena. Section 5 presents our experiment results and analysis. Section 6 provides further discussion of our method, results and implications for future directions.
2 Emergence and Self-organization
“Emergence refers to the arising of novel and coherent structures, patterns, and proerties during the process of self-organization” (Goldstein, 1999). The emergence can be understood as the process by which new order or pattern forms - “something new appears” - in the evolution of a collective system that cannot be described by its dynamical equations, and which can be exploited by this system through various mechanisms to gain additional functionalities or capabilities (Crutchfield, 1994). In the context of LLMs, the emergence can be understood as the complex behavior by which new learning capabilities are acquired by the LLM from the collective interactions among a multitude of simple elements (i.e., the neurons which follow simple and consistent rules within the neural network). As the scale of the LLM increases, these interactions can become increasingly intricate, contributing in a synergistic way to new patterns; thus, the whole system significantly surpasses the sum of its individual parts in terms of functionality.
Self-organization is the process of pattern formation through internal interactions (Correia, 2006; Ay et al., 2007). Self-organization refers to the ability of a collective system to drive itself towards a more ordered state and maintain its functionality in the face of perturbations. Just as in natural systems where intricate patterns and behavior emerge from the interactions of individual elements following basic rules, self-organization in the context of neural networks implies the development of intricate structures and capabilities as a result of the interactions between neurons, akin to Turing Patterns where complex patterns emerge from simple rules in natural systems (see Appendix A.3.2). In contrast to prior work on observing independent LLMs at different scales, we focus on how untrained models, during training, gradually manifest increasing levels of order over multiple scales. Understanding and quantifying the degree of order-ness appearing or developing during the training of LLMs requires a strategy of estimating the rules from small data through geometric arguments (i.e., only weights of neuronal interactions are known). Towards this end, the multifractal theory (Evertsz & Mandelbrot, 1992; Salat et al., 2017; Song et al., 2005) allows us to identify rules and patterns across multiple scales.
@@ 3 Related Work
While the optimization loss decreases with increased model size, the task performance exhibits complex behavior, showing significant improvement only after a model exceeds a threshold in size, a phenomenon known as emergent abilities (EAs) (Wei et al., 2022a). Recently, the emergence has gained increasing attention in LLMs, first within the technique of ‘chain-of-thought’ on the performance of solving complex mathematical problems (Wei et al., 2022b). According to (Schaeffer et al., 2023), the scope of ‘emergence’ is defined as: 1). Large model Only: abilities that can not be found in small models; 2). Sudden change on monitored metrics: the model’s behavior has a sudden change, which means it cannot be inferred from the small model’s performance.
In emergence interpretbility area  (Fu et al., 2023),  Hu et al. introduced a quantitative evaluation strategy (i.e., PASSUNTIL), which uses massive sampling to achieve theoretically infinite resolution. Derived from the quantization hypothesis,  Michaud et al. presented the quantization model for neural scaling laws, which aims to explain the observed power-law decrease in loss concerning model and data size, as well as the phenomenon of EAs with scale.  Gurnee et al. investigated the concept of universal neurons in GPT-2 language models, finding that a small percentage of neurons are universal, exhibiting consistent behavior across models trained with different random seeds. In contrast,  Schaeffer et al. challenged the notion of EAs in LLMs and proposed that the appearance of emergent abilities is influenced by the choice of evaluation metrics rather than actual changes in model behavior.
Furthermore, LLMs (e.g., ChatGPT, LLaMA) encounter limitations in domain-specific tasks. They often lack depth and accuracy in specialized areas, and exhibit a decrease in general capabilities when fine-tuned, particularly analysis ability in small sized models (Gaur & Saunshi, 2023; Zheng et al., 2024; Chan et al., 2023). According to Wei et al. (2023), smaller LLM relies primarily on semantic priors from pretraining when presented with flipped in-context labels, whereas larger models can override priors, demonstrating reduced reliance on semantic priors even with potentially stronger priors due to larger scale. Jaiswal et al. proposed the existence of essential sparsity when directly removing weights with the smallest magnitudes in one-shot without re-training. While prior work observed emergent outcomes, the complex internal dynamics and evolving interactions during training remain overlooked. In this work, we argue that understanding emergence requires investigating the neuronal-level processes whereby self-organization arises from initially random interactions as training progresses.
@@ 4 Methodology
4.1 LLMs’ Neurons as Interacting Agents in a Network
To illustrate our method, let us define the components of a neural network (NN). Let $N$ denote a neural network. $N$ is composed of several blocks $B_{1},B_{2},\cdots,B_{k}$. Each block $B_{i}$ contains multiple layers $L_{1}^{i},L_{2}^{i},\cdots,L_{m_{i}}^{i}$. Each layer $L_{j}^{i}$ comprises multiple neurons $n_{1}^{ij},n_{2}^{ij},\cdots,n_{p_{ij}}^{ij}$. Neurons between layers are interconnected by weights $w_{ab}\in W$, where $a$ and $b$ denote the indices of the neurons. There are no connections between neurons within the same layer.
Figure 3: The transformation from a neural network to a neural interaction network.
Figure 4: (a) shows examples of our weighted box covering method, illustrating two neurons randomly selected from the NIN of pythia 1B model at the $143,000^{th}$ epoch. The radius of each box, $r$, is determined by the distance to its neighboring neurons, with $N$ representing the number of neurons contained within a box. (b) demonstrates the power-law relationship between $N$ and $r$.
The structure of a NN provides information about its architecture and connections. Consequently, our proposed multifractal analysis investigates the interactions among neurons, with a specific emphasis on identifying the shortest paths or distances. In NNs, larger weights typically indicate stronger interactions, suggesting shorter distance between two neurons in our network analysis. This motivates us to transform the NN into a so called “Neuron Interaction Network”, defined as follows.
@@ 4.1 LLMs’ Neurons as Interacting Agents in a Network
NIN is a directed weighted graph $G=(V,E)$ with a weight transfer function $f$, where:
[1]. $v\in V$ denotes the set of neurons in the NN; since nodes in the NIN are the same as those in the NN, we do not differentiate between them in this article;
[2]. $e_{a,b}\in E,(a,b\in V)$ represents a link between original neurons, with a new weight $d_{ab}$ determined by $f$;
[3]. $f:\mathcal{R}\rightarrow\mathcal{R}$ decides the edge distance in the NIN, being calculated based on original weights among neurons;
[4]. Blocks $B$ and layers $L$ are preserved by the NIN structure.
We assume that the function $f$ has the property of Absolute Value and Decreasing, where the absolute value escapes the influence of the negative output on weight calculation, and decreasing indicates we only target the shortest path for analysis. Thus, we formulate $f$ as follows:
$f(w_{ij})=\begin{cases}\frac{1}{|w_{ij}|},&\text{if }w_{ij}>\epsilon\\ 0,&\text{otherwise}\end{cases},\forall i,j\in V$ (1)
where $\epsilon$ is the threshold of weights, $w_{ij}$, for this quantifiable network model to guarantee the sparsity. We also explore different functions $f$ and provide the ablation studies in section Section 5.4.
Since LLMs contain a huge number of neurons, it is inefficient to do the analysis on the whole NIN. To this end, we define a Sampled Neural Interaction Network to improve the analysis efficiency.
@@
A SNIN represents a sampled subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ of the original NIN. In each layer $L_{i}$, we sample a subset of nodes $V_{i}^{\prime}\in L_{i}$, and add them to SNIN and having weights $E^{\prime}=\{e_{ab}|e\in E,a,b\in V^{\prime}\}$.
To enhance the precision and efficiency on NIN analysis, we adopt a strategy where we average the outcomes from analyses of 10 randomly selected SNIN sets in NIN. As substantiated in Section A.1.2, this sampling methodology does not alter the estimation of the fractal dimensions. For clarity and simplicity, we consistently use the notation $G=(V,E)$ and refer to the NIN instead of SNIN in subsequent sections.
4.2 Neuron-based Multifractal Analysis (NeuroMFA)
To perform multifractal analysis of LLMs, we extend the box-covering and box-growing methods (Song et al., 2005; Xiao et al., 2021) to capture the local (based on each neuron) fractal properties of NINs. Specifically, let us consider neuron $v_{i}$ as a box with an initial radius of zero at layer $l$ and indexed by $i$. We then increase the radius of the box and record the mass distribution (i.e., the number of nodes covered by the box). First, we define $N_{l,i}(r)$ as the number of neurons connected to neuron $v_{l,i}$ within a radius $r$. For a specific neuron $v_{i}$ at layer $l$ and radius $r$, $N_{l,i}(r)$ is calculated as follows:
$N_{l,i}(r)=\sum_{v_{j}\in L_{l+1}}\mathbf{1}_{\{d_{ij}\leq r\}},$ (2)
where, $\mathbf{1}_{\{d_{ij}\leq r\}}$ is an indicator function that takes the value 1 when the distance between neuron $v_{i}$ and $v_{j}$, $d_{ij}$, is less than or equal to $r$, and 0 otherwise. Thus, $N_{l,i}(r)$ computes the number of neurons in the next layer(s) connected to neuron $v_{i}$ within the radius $r$.
The process continues with increasing the box radius until all the neighbors of the neuron are covered. The method captures the log-log relationship between the box radius $r$ and the mass distribution $N_{l,i}(r)$ in the box, and indicates the fractal properties of the neuron interactions (see Figure 4 and Table 2, Section C.1.4).
Observation 4.1.
In a NIN, the relation between the mass distribution $N(r)$ and box radius $r$ can be expressed as:

### v5

# Neuron-based Multifractal Analysis of Neuron Interaction Dynamics in Large Models

@@ Abstract
In recent years, there has been increasing attention on the capabilities of large models, particularly in handling complex tasks that small-scale models are unable to perform. Notably, large language models (LLMs) have demonstrated “intelligent” abilities such as complex reasoning and abstract language comprehension, reflecting cognitive-like behaviors. However, current research on emergent abilities in large models predominantly focuses on the relationship between model performance and size, leaving a significant gap in the systematic quantitative analysis of the internal structures and mechanisms driving these emergent abilities. Drawing inspiration from neuroscience research on brain network structure and self-organization, we propose (i) a general network representation of large models, (ii) a new analytical framework—Neuron-based Multifractal Analysis (NeuroMFA)-for structural analysis, and (iii) a novel structure-based metric as a proxy for emergent abilities of large models. By linking structural features to the capabilities of large models, NeuroMFA provides a quantitative framework for analyzing emergent phenomena in large models. Our experiments show that the proposed method yields a comprehensive measure of network’s evolving heterogeneity and organization, offering theoretical foundations and a new perspective for investigating emergent abilities in large models.
11footnotetext: * Correspondence to: Xiongye Xiao <xiongyex@usc.edu>, Paul Bogdan <pbogdan@usc.edu>
@@ 1 Introduction
The advent of foundation models triggered a paradigm shift in artificial intelligence (AI), demonstrating unparalleled capabilities across many domains, particularly natural language processing  (Chang et al., 2024; Brown et al., 2020; Chowdhery et al., 2023; He et al., 2020). Recent studies observed that, compared to traditional small-scale neural networks (NNs), large language models (LLMs) exhibit advanced cognitive and reasoning abilities and a higher level of comprehension  (Brown et al., 2020; Bubeck et al., 2023; Achiam et al., 2023). This behavior, where larger models develop qualitatively new abilities absent in smaller ones, is referred to as "emergent abilities" (Srivastava et al., 2022; Wei et al., 2022a). Recent works have focused on using scaling laws and other analytical methods to study the relationship between model size and the development of these new capabilities, suggesting that emergent abilities may arise predictably as models scale up (Kaplan et al., 2020; Hu et al., 2023a; b; Michaud et al., 2023; Schaeffer et al., 2024).
Figure 1: (a) The proposed paradigm shift for assessing the degree of emergence in large models. (b) Comparison of traditional methods (top) and our approach (down); we study the emergent abilities of large models by analyzing structural dynamics during training.
Existing research, limitations and motivation. As depicted in Fig. 1, previous research on the "emergence" in large models and particularly LLMs (Wei et al., 2022a; Fu et al., 2023; Zhong et al., 2024; Schaeffer et al., 2024) focuses on the observation-based analysis and the model size (e.g., number of parameters, model depth). Traditional metrics (Paperno et al., 2016; Sakaguchi et al., 2021; Bisk et al., 2020) have largely assessed LLMs’ emergent abilities by capturing the relationship between model’s size and output performance. However, these approachs have several limitations: (i) observation-based methods often require exhaustively testing various outputs, making it challenging to establish a unified standard for evaluating the full scope of a large model’s emergent abilities; (ii) the approaches treat the model as a black box, limiting the understanding of the internal mechanisms that drive the model’s behavior and emergent abilities; (iii) relying solely on model size to measure performance overlooks other critical factors that significantly influence the model’s capabilities, such as architecture design, data quality and training dynamics. To address these limitations, our approach emphasizes structure-based analysis by examining the dynamic changes in a model’s internal structure during the training process. The iterative training process is crucial for the improvement and development of capabilities in large models, as it enables the model to transition from having no abilities to exhibiting emergent abilities, as illustrated in Fig. 2. Current work has demonstrated that large models undergo significant structural changes in their internal mechanisms and dynamic characteristics throughout training (Ding et al., 2023). Moreover, inspired by neuroscience research on how the brain’s network microstructure supports its functions (Bullmore & Sporns, 2009; Bassett & Sporns, 2017), we aim to develop a similar framework for the structural analysis of large models. This framework seeks to link the internal architecture and interaction patterns of large models with their emergent capabilities, providing deeper insights into the mechanisms that drive their complex behaviors. By adopting a structure-based analysis approach, we can move beyond treating models as black boxes and gain a deeper understanding of how internal structural dynamics connect to the development of advanced functionalities in large models.
@@
Figure 2: Improvement of Pythia responses with increasing training steps, as indicated by the progression from incoherent to coherent answers across diverse categories (blue for questions with standard answers, red for semi-open-ended questions, and yellow for open-ended questions).
Our approach and its connection to neuroscience. To address the above limitations, our research bridges the existing gap by analyzing the internal network structures of large models and linking the structure-based properties to the models’ emergent capabilities. In the field of neuroscience, extensive research has established that the brain’s network microstructure is fundamental to its functionality, with model organisms such as Drosophila and mice serving as key examples (Mizutani et al., 2013; Shih et al., 2020; Xiao et al., 2021; Yin et al., 2020; Coletta et al., 2020). These studies emphasize the critical role of network structure in supporting and determining functional capabilities, inspiring us to explore the relationship between the structure and the emergent abilities of large artificial models. In the study of human intelligence, the brain is considered a self-organizing complex system (Kelso, 1995; Singer, 1986; Pribram, 2018; Antonello et al., 2022). The manifestation of intelligence and human behavior—from neurons to mind—is governed by the generic process of self-organization. A key aspect of studying intelligence involves quantitatively analyzing and investigating the complex interactions among numerous neurons (Sporns et al., 2004; Bassett & Sporns, 2017). Self-organization phenomena, widely observed in natural systems, involve numerous micro-level interactions leading to complex macro-level behaviors, and the emergence can be understood by modeling and analyzing the self-organizing micro-structure (Turing, 1990) (see Appendix A.3). Inspired by this, we aim to explore whether the emergent behaviors exhibited by large models can be understood and connected through structural self-organization. By abstracting large models as network representation, we can analyze their interaction structures and correlate them with capabilities. Drawing an analogy to structural self-organization, we propose relevant structure-based metrics that enable us to quantify the degree of emergence (this is a structure-based metric that can be viewed as a proxy for the emergent abilities of large models) in large models.
Our contributions. In summary, the contributions of our work are three-fold:
(i) General network representation of LLMs: We propose a unified network representation for large models, where large models are represented as Neuron Interaction Networks (NINs), enabling the structural analysis of artificial neuron interactions.
(ii) Structural analysis framework: We introduce neuron-based multifractal analysis (NeuroMFA), a framework that quantifies the regularity and heterogeneity of NINs by statistically analyzing the fractal properties of neuron interactions. This framework measures the neuron interaction dynamics at the micro-level and it is the first flexible framework that seamlessly integrates techniques for self-organization, fractal and structural analysis to estimate the degree of emergence of large models.
(iii) Structural Emergence Metric: We propose a novel metric based on NeuroMFA to quan tify the degree of structural emergence in large models by analyzing the interaction dynamics of neurons during the training process, opening a new direction for the study of "emergence" in large models.
@@ 2 Related work
The concept of emergent ability has gained prominence in LLM research, initially through the ’chain-of-thought’ technique for complex mathematical problem-solving (Wei et al., 2022b). The interpretability and quantification of emergent ability were advanced with the PASSUNTIL evaluation strategy (Fu et al., 2023), utilizing massive sampling for high-resolution analysis (Hu et al., 2023b).  Michaud et al. (2023) proposed a quantization model to elucidate neural scaling laws, addressing both the power-law decrease in loss with increasing model and dataset sizes, and the development of new abilities as models scale up. However,  Schaeffer et al. (2024)questioned the notion of emergent abilities in LLMs, arguing that the appearance of these abilities may be more a consequence of the evaluation metrics chosen, rather than fundamental changes in the model’s behavior. This critique has sparked further inquiry into how evaluation strategies influence our understanding of emergent abilities in large models.
Figure 3: The transformation from a partial neural network (NN) in large models into a neuron interaction network (NIN).
Following this,  Chen et al. (2024) proposed a quantitative approach to measuring "emergence" by comparing entropy reductions at semantic and token levels.  Gurnee et al. (2024) explored the concept of "universal neurons" in GPT-2 models, discovering that a small subset of neurons exhibits consistent behavior across models trained with different random seeds, suggesting a neuron-level mechanism that may drive emergent phenomena. Most recently,  Du et al. (2024) introduced a new perspective, emphasizing pre-training loss as a critical driver of emergent ability in large models. Their study showed that models with lower pre-training losses exhibit emergent abilities, independent of model size or dataset scale.
@@ 3 Methodology
In this section, we introduce the motivation behind our approach and outline its three main components. Briefly, we first represent a large model as a neuron interaction network. Then, we propose the neuron-based multifractal analysis (NeuroMFA). Finally, based on NeuroMFA, we propose a new metric as a proxy for emergent abilities of large models by quantifying and analyzing the dynamic changes in their structure.
3.1 Motivation: Connecting self-organization to emergence
Self-organization is a foundational concept for understanding emergent phenomena in natural systems (Goldstein, 1999; Crutchfield, 1994). It describes how systems gradually evolve through local interactions to develop more diverse interaction patterns and achieve overall order, ultimately leading to emergence (Correia, 2006; Ay et al., 2007). This process involves two key aspects: the development of new interaction patterns (reflecting increased heterogeneity in interactions) and the formation of orderly structures (representing enhanced regularity in global structure). Emergence, in this context, refers to the novel macro-level properties formed during the process of self-organization—characteristics that arise through collective interactions and cannot be fully explained by the dynamics of individual components. Inspired by these principles, we extend the concept of self-organization to study degree of emergence in large artificial models (please refer to Appendix A.3.3) and design network-based metrics to quantify the two key aspects of self-organization: heterogeneity (capturing the diversity of neuron interactions) and regularity (reflecting the order in global network structure). By introducing neuron-based multifractal analysis (NeuroMFA), we provide a unified framework to investigate these dimensions, enabling a quantitative analysis of the relationship between structure and emergent capabilities in LLMs.
3.2 Neuron interaction network
To illustrate our approach, we begin by defining the structure of a neural network (NN) as observed within large models. A large model is usually composed of multiple NN layers, represented as $L_{1},L_{2},\cdots,L_{k}$, where $k\in\mathbb{N}^{+}$ denotes the number of layers. Within each layer $L_{j}$, the neurons are specified as $n_{1}^{j},n_{2}^{j},\cdots,n_{p_{j}}^{j}$, where $p_{j}\in\mathbb{N^{+}}$ denotes the number of neurons in layer $L_{j}$. Neurons between layers are interconnected by weights $w_{ab}\in W$, where $a$ and $b$ denote the indices of the neurons. For instance, $a$ could represent $n_{1}^{j}$ and $b$ could represent $n_{1}^{j+1}$. Notably, neurons in one layer connect only to neurons in the subsequent layer.
[…]
The structure of an NN contains key information about its architecture and neuron connections. However, applying network and structure analysis to NNs presents challenge: (i) the presence of negative weights can create inconsistencies in distance-based metrics, complicating the interpretation of interactions; (ii) defining measures of distance and shortest path is crucial for studying network structures, where stronger connections typically imply shorter distances (Newman, 2001). To address these issues, we first transform NNs into the Neuron Interaction Networks (NINs), defined as follows.
@@ 3.2 Neuron interaction network
NIN is a weighted directed graph $G=(V,E)$, where:
[1] Node set: $V$ denotes the set of neurons (nodes) in the NIN, directly inherited from the NN;
[2] Edge set: $E\subseteq V\times V$ represents the directed edges between nodes, where each edge $e_{a,b}\in E$ connects nodes $a$ and $b$. Each edge has a non-negative weight $\omega_{ab}=|w_{ab}|^{p}$, where $w_{ab}$ is the original weight in the NN, $|w_{ab}|$ reflects the interaction strength between neurons, and $p=-1$ is a parameter that inversely maps the original weight to the edge distance (Newman, 2001; Liu et al., 2017).
[3] Shortest path distance: the distance $d_{ij}\in\mathbb{R}_{0}^{+}$ between two nodes $i$ and $j$ is defined as the shortest path distance, calculated as the minimum sum of edge weights over all possible paths between the nodes. A path $P(i,j)=(v_{i},v_{1},v_{2},\cdots,v_{j})$ consists of a valid sequence of neurons and edges connecting $i$ to $j$, spanning multiple layers. The formal expression for the distance is given by:
$d_{ij}=\min_{P(i,j)}\left(\sum_{(u,v)\in P(i,j)}\omega_{uv}\right)+\lambda\cdot|P(i,j)|^{\gamma}.$ (1)
The term $|P(i,j)|$ is defined as the number of edges in the path $P(i,j)$ connecting nodes $i$ and $j$, representing the length of the shortest path. This distance metric accumulates the edge weights $\omega_{uv}$ along each valid path and incorporates a weighted term based on the path length. The parameters $\lambda\in\mathbb{R}$ and $\gamma\in\mathbb{R}$ balance the influence of edge weights and path lengths, respectively, allowing for a comprehensive measure of distance that accounts for both interaction strengths and the complexity of the connectivity paths.
[4] Neighbors $\mathcal{N}(v_{i})$: To facilitate the proposed analysis of neuron-based local interaction patterns in the NIN, we define the neighbors of a neuron $v_{i}\in V$ as the set of neurons whose shortest path distance from $v_{i}$ is less than or equal to a predefined threshold distance $d_{\text{threshold}}$. Formally, the neighbor set $\mathcal{N}(v_{i})$ is defined as:
$\mathcal{N}(v_{i})=\{v_{j}\in V\mid d_{ij}\leq d_{\text{threshold}}\},$ (2)
where $d_{ij}$ denotes the shortest path distance between neurons $v_{i}$ and $v_{j}$. This definition identifies all neurons that are sufficiently close to $v_{i}$ in terms of network proximity, enabling us to efficiently and systematically explore local node-based interaction dynamics.
In the subsequent NeuroMFA framework, these neighborhood relationships are leveraged to conduct multifractal analysis, revealing multi-scale structures and complex interaction patterns within the network. Since LLMs contain a huge number of neurons, it is inefficient to apply the fractal analysis to the whole NIN. To this end, we define a Sampled Neural Interaction Network to improve the analysis efficiency and the generalization capabilities of our method.
@@
An SNIN represents a sampled subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ of the original NIN. In each layer $L_{i}$, we sample a subset of nodes $V_{i}^{\prime}\in L_{i}$ and add them to SNIN with weights $E^{\prime}=\{e_{ab}|e\in E,a,b\in V^{\prime}\}$.
To enhance the precision and efficiency of the NIN analysis, we adopt a strategy where we average the outcomes from the study of 10 randomly selected SNIN sets in NIN. As substantiated in Section A.1.3, this sampling methodology does not alter the estimation of the fractal dimensions. For clarity and simplicity, we consistently use the notation $G=(V,E)$ and refer to the NIN instead of SNIN in subsequent sections.
3.3 Neuron-based multifractal analysis
To perform multifractal analysis of LLMs, we extend the box-covering and box-growing methods (Song et al., 2005; Evertsz & Mandelbrot, 1992; Salat et al., 2017; Xiao et al., 2021) to capture the local neuron-based fractal properties of NINs. Specifically, let us consider a neuron $v_{i}$ as a box with an initial radius of zero at layer $l$, indexed by $i$. We then increase the radius of the box and record the mass distribution $N_{l,i}(r)\in\mathbb{N}^{+}$ (i.e., the number of nodes covered by the box), as shown in Fig. 4 (a). That is to say, $N_{l,i}(r)$ denotes the number of neighbors of neuron $v_{l,i}$ within a radius $r$.
For a specific neuron $v_{i}$ at layer $l$ and radius $r$, $N_{l,i}(r)$ is calculated as $N_{l,i}(r)=\sum_{v_{j}\in\mathcal{N}(v_{i})}\mathbf{1}_{\{d_{ij}\leq r\}}$, where $\mathbf{1}_{\{d_{ij}\leq r\}}$ is an indicator function that takes the value 1 when the distance $d_{ij}$ between neighbor neurons $v_{i}$ and $v_{j}$ is less than or equal to $r$ and 0 otherwise. The process continues to increase the box radius until all the neighbors of the neuron are covered. Based on this approach, we capture the log-log relationship between the box radius $r$ and the mass distribution $N_{l,i}(r)$ in the box as shown in Fig. 4 (b), indicating the fractal properties of the neuron interactions. The log-log relationship between $N(r)$ and $r$ through the whole NIN presented in Appendix Table 9 leads to the following observation.
Observation 1 (Fractal Properties).
In an NIN, the relation between the mass distribution $N(r)$ and box radius $r$ can be expressed as: